If i bought an item on sale for 266 dollars after receiving 65% off, what is the original price of that item?
Answer by jim_thompson5910(35256)

Let x = original price


(Original Price) - (65% of Original Price) = Sale Price

(x) - (65% of x) = 266

x - 0.65x = 266

0.35x = 266

x = 266/0.35

x = 760


The original price of the item is $760